OpenStudy (anonymous):

Doing series convergent tests, and on a roots test, I get to a limit that I don't know how to prove. Prove: Limit n to infinity of ((n!)^5)/((5n)!) = 0 Do i need to do L'Hospitals rule? If so, how do I take the derivative of the top and bottom.

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} \frac{ n!^5 }{ (5n)! }\]

OpenStudy (anonymous):

How did you come about this limit?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{ n!^{5n} }{ (5n)!^{n} }\] This is the origional sum, when doing the roots test, you factor out n, take the nth root, and take the limit of the absolute value of a sub n.

OpenStudy (anonymous):

I would say that this series does not converge.

OpenStudy (da_scienceman):

Did u try d'Alemberts ratioj test?